# OmniGuard: An Efficient Approach for AI Safety Moderation Across Modalities

### v1

@@ Abstract
The emerging capabilities of large language models (LLMs) have sparked concerns about their immediate potential for harmful misuse. The core approach to mitigate these concerns is the detection of harmful queries to the model. Current detection approaches are fallible, and are particularly susceptible to attacks that exploit mismatched generalization of model capabilities (e.g., prompts in low-resource languages or prompts provided in non-text modalities such as image and audio). To tackle this challenge, we propose OmniGuard, an approach for detecting harmful prompts across languages and modalities. Our approach (i) identifies internal representations of an LLM/MLLM that are aligned across languages or modalities and then (ii) uses them to build a language-agnostic or modality-agnostic classifier for detecting harmful prompts. OmniGuard improves harmful prompt classification accuracy by 11.57% over the strongest baseline in a multilingual setting, by 20.44% for image-based prompts, and sets a new SOTA for audio-based prompts. By repurposing embeddings computed during generation, OmniGuard is also very efficient (120$\approx$ 120$\times$ 120 × faster than the next fastest baseline). Code and data are available at https://github.com/vsahil/OmniGuard.
@@
for AI Safety Moderation Across Modalities
@@ 1 Introduction
The rapid rise of capabilities in large language models (LLMs) has created an urgent need for safeguards to prevent their immediate harmful misuse as they are deployed to human users en masse Bommasani et al. (2022). Moreover, these safeguards are critical for defending against future potential harms from LLMs Bengio et al. (2024). Standard safeguard approaches broadly include approaches such as safety training using reinforcement learning from human feedback (Ouyang et al., 2022a; Leike et al., 2018) or using pre-trained guard models that classify the safety of an input prompt  (OpenAI, 2025; Inan et al., 2023; Han et al., 2024).
With these safeguards in place, harmful prompts in high-resource languages, e.g., English, are successfully detected. However, harmful prompts in low-resource languages can often bypass these safeguards (Deng et al., 2024; Yong et al., 2024; Yang et al., 2024), i.e., jailbreaking the LLM. Modern LLMs are vulnerable to attacks not only from low-resource natural languages, but also from artificial cipher languages, e.g., base64 or caesar encoding of English prompts (Wei et al., 2023; Yuan et al., 2024a). This phenomenon also extends beyond text to jailbreaking multimodal LLMs (MLLMs) using modalities such as images (Gong et al., 2025; Liu et al., 2024b) or audio (Yang et al., 2025).
Wei et al. (2023) argue that these attacks are successful due to mismatched generalization, a scenario in which the model’s safety training does not generalize to other settings, but general performance does. This might happen since pretraining data often includes more diverse data than that available for safety finetuning (Ghosh et al., 2024b). In this work, we defend against attacks that exploit the mismatched generalization of the safety training of LLMs and MLLMs. Specifically, we defend against attacks that utilize low-resource languages, both natural and cipher languages, as well as attacks employing other modalities, such as images and audio.
@@
Compared to popular guard models such as LlamaGuard (Inan et al., 2023), AegisGuard (Ghosh et al., 2024a), or WildGuard (Han et al., 2024) among others, OmniGuard does not require training a separate LLM specifically to detect harmfulness. By building a classifier that uses the internal representations of the main LLM or MLLM, OmniGuard completely avoids the overhead of passing the prompt through a separate guard model, making it very efficient.
@@ 2 Methodology
OmniGuard seeks to robustly detect harmful prompts, regardless of their language or modality. We first leverage the tendency of LLMs and MLLMs to create universal representations that are similar across languages (Wendler et al., 2024; Zhao et al., 2024) and across modalities (Wu et al., 2024; Zhuang et al., 2024) in Section 2.1, and then use them to train harmfulness classifiers that robustly detect harmful inputs in Section 2.2.
@@ 2.1 Finding language-agnostic representations in an LLM
This procedure can be generalized to new different modalities rather than different languages by changing which embeddings are being used. For example, to determine if internal representations of an MLLM are aligned across modalities, we replace embeddings for a translated piece of text with embeddings from a different modality (e.g. a text caption and its corresponding image, or a text transcription and its corresponding audio clip). See experimental details in Section 3.
@@ 3 Experimental Setup
Table 1 and Table 2 give details on all the models and datasets for this section.
@@
General Flores200 Team et al. (2022) Muennighoff/flores200 997
MM-Vet v2 Yu et al. (2024b) whyu/mm-vet-v2 517
SST-2 (Socher et al., 2013) stanfordnlp/sst2 1000
Text Aegis AI Content Safety Dataset Ghosh et al. (2024b) nvidia/Aegis-AI-Content-Safety-Dataset-1.0 10,800
MultiJail Deng et al. (2024) DAMO-NLP-SG/MultiJail 315
Xsafety Wang et al. (2024a) ToxicityPrompts/XSafety 28,000
RTP-LX de Wynter et al. (2025) ToxicityPrompts/RTP-LX 30,300
AyaRedTeaming Aakanksha et al. (2024) CohereLabs/aya_redteaming 2662
Thai Toxicity tweets Sirihattasak et al. (2018) tmu-nlp/thai_toxicity_tweet 3,300
Ukr Toxicity Dementieva et al. (2024) ukr-detect/ukr-toxicity-dataset 5,000
HarmBench (HB) (Mazeika et al., 2024) walledai/HarmBench 400
Forbidden Questions (FQ) (Shen et al., 2024a) TrustAIRLab/forbidden_question_set 390
Simple Safety Tests (Vidgen et al., 2024) walledai/SimpleSafetyTests 100
SaladBench (SaladB) (Li et al., 2024a) walledai/SaladBench 26,500
Toxicity Jigsaw (TJS) cjadams et al. (2017) Arsive/toxicity_classification_jigsaw 26,000
Toxic Text (Corrêa, 2023) nicholasKluge/toxic-text 41,800
AdvBench (Zou et al., 2023a) walledai/AdvBench 520
CodeAttack (Ren et al., 2024) https://github.com/AI45Lab/CodeAttack 3120
Vision JailBreakV-28K (Luo et al., 2024) JailbreakV-28K/JailBreakV-28k 8,000
VLSafe (Chen et al., 2024c) YangyiYY/LVLM_NLF 1,110
FigStep (Gong et al., 2025) https://github.com/wangyu-ovo/MML 500
MML SafeBench (Wang et al., 2024b) https://github.com/wangyu-ovo/MML 2,510
Hades (Li et al., 2024e) Monosail/HADES 750
SafeBench (Ying et al., 2024) Zonghao2025/safebench 2,300
MM SafetyBench (Liu et al., 2024b) PKU-Alignment/MM-SafetyBench 1680
RedTeamVLM (Li et al., 2024b) MMInstruction/RedTeamingVLM 200
VLSBench (Hu et al., 2025) Foreshhh/vlsbench 2,240
Audio VoiceBench (Alpacaeval) Chen et al. (2024d) hlt-lab/voicebench 636
AIAH (Yang et al., 2025) https://github.com/YangHao97/RedteamAudioLMMs 350
@@
Model name Citation HuggingFace ID Rough Parameter Count
General Llama3.3-70B-Instruct (Grattafiori et al., 2024) meta-llama/Llama-3.3-70B-Instruct 70B
Molmo-7B (Deitke et al., 2024) allenai/Molmo-7B-D-0924 7B
LLaMA-Omni 8B (Fang et al., 2025) ICTNLP/Llama-3.1-8B-Omni 8B
Kokoro (Hexgrad, 2025) hexgrad/Kokoro-82M 82M
Text LlamaGuard 1 Inan et al. (2023) meta-llama/LlamaGuard-7b 7B
LlamaGuard 2 Inan et al. (2023) meta-llama/Meta-Llama-Guard-2-8B 8B
LlamaGuard 3 Inan et al. (2023) meta-llama/Llama-Guard-3-8B 8B
AegisGuard Permissive Ghosh et al. (2024a) nvidia/Aegis-AI-Content-Safety-LlamaGuard-Permissive-1.0 7B
AegisGuard Defensive Ghosh et al. (2024a) nvidia/Aegis-AI-Content-Safety-LlamaGuard-Defensive-1.0 7B
WildGuard Han et al. (2024) allenai/wildguard 7B
HarmBench (mistral) Mazeika et al. (2024) cais/HarmBench-Mistral-7b-val-cls 7B
HarmBench (llama) Mazeika et al. (2024) cais/HarmBench-Llama-2-13b-cls 13B
DuoGuard Deng et al. (2025) DuoGuard/DuoGuard-1B-Llama-3.2-transfer 1B
PolyGuard Kumar et al. (2025) ToxicityPrompts/PolyGuard-Qwen 7B
Vision Llama Guard 3 Vision (Chi et al., 2024) meta-llama/Llama-Guard-3-11B-Vision 11B
VLMGuard (Du et al., 2024) --- 2.2M
LLavaGuard (Helff et al., 2025) AIML-TUDA/LlavaGuard-7B-hf 7B
Table 2: Model and baseline details.
@@ 3.1 Selecting universal layers via the U-Score
To select language-agnostic layers, we use a dataset of translated sentence pairs spanning various languages. Specifically, we use sentences in 53 natural languages from the Flores200 dataset and additionally translate the sentences into 20 cipher languages (using encodings such as Caesar shifts, base64, hexadecimal); see a full list in Appendix A. We extract embeddings from each layer of Llama3.3-70B-Instruct for the sentences in all 73 languages and use them to compute the U-Score (averaged over languages). Fig. 2 shows the U-Score as a function of layer depth. For natural languages (blue curve), the U-Score peaks in the middle layers of the model, with the highest values in layer 57 (out of 81 layers). For cipher languages (red curve), the U-Score is much lower than for natural languages, suggesting the model fails to represent semantic similarity in these languages (see analysis in Section 5).
@@
To select layers aligned between images and captions, we use the MM-Vet v2 dataset, a popular dataset for MLLM evaluation containing 517 examples, each consisting of a text question paired with one or more images. We generate captions for each image using a captioning model (Molmo-7B) and then extract embeddings for each image and its corresponding caption using an MLLM (also Molmo-7B) and use them to compute the U-Score, which peaks in layer 22 (out of 28 layers; see Fig. 2 green curve).
To select layers aligned between text and audio, we use the audio version of the Alpacaeval dataset from VoiceBench, a dataset of 636 audio-transcript pairs. We extract embeddings from each layer of an MLLM (LLaMA-Omni 8B) and use them to compute the U-Score, which peaks at layer 21 (out of 32 layers; see Fig. 2 purple curve).
@@ 3.2.1 Setup for multilingual text attacks
Following Section 3.1, we build a classifier that takes as input embeddings from layer 57 of Llama3.3-70B-Instruct. As training data, we randomly select 2,800 examples from the Aegis AI Content Safety dataset, balancing the benign and harmful classes. Notably, this dataset is about 18$\times$ smaller than the training data used by our baseline methods. We translate these English examples to 52 other natural languages (via the Google Translate API) and 20 cipher languages (using fixed rules), totaling 73 languages. We train OmniGuard using only half the languages (see list in Appendix A).
@@
We compare to many popular guard models (see Table 2) middle row. Notably, DuoGuard and PolyGuard were trained to detect harmful prompts across multiple languages. For a more direct comparison, we also compare to finetuned versions of DuoGuard and PolyGuard using the same 37 languages we use to train OmniGuard; Following the original PolyGuard paper, we finetuned these models using LoRA (Hu et al., 2021) for all linear layers with rank 8 and alpha 16 for one epoch with a learning rate of $42e-$ 4.
@@
We evaluate on several common text attack benchmarks (see Table 1). We additionally evaluate on three benchmarks from CodeAttacks that transform a harmful query as a list, a stack, or as a string in a Python program, obfuscating the harmfulness. For evaluation in this setup, we transform the harmful prompts from AdvBench and benign prompts from Toxicity Jigsaw datasets in the three code formats and subsampl the Toxicity Jigsaw dataset to be of the same size as Advbench. Note that, for this experiment, we only trained OmniGuard on the English subset of the training dataset.
@@ 3.2.2 Setup for vision attacks
Following Section 3.1, we build a classifier that takes as input embeddings from layer 22 of Molmo-7B. As training data, we use 2000 image-query pairs randomly sampled from the JailBreakV-28K dataset and 1024 image-query pairs sampled from the VLSafe dataset as the harmful datapoints and 517 image-query from the MM-Vet v2 dataset as the benign datapoints.
@@
We compare to guard models that take an image or image-text pair and output a binary harmfulness classification (see Table 2 bottom row). We train VLMGuard on the same training data as OmniGuard.
@@
We evaluate detecting image/text attacks using several datasets (see Table 1). FigStep and MML Safebench are typographic attacks that embed a harmful prompt in an image. MML Safebench further encrypts a harmful prompt in several variants, such as rotation, mirror images, word replacement, and with base64 encoding. Hades and Safebench consist of images and text queries where the text itself is harmful. MM-safetybench, RTVLM, and VLSBench consist of an image and a query where the text query is seemingly benign, but when combined with the respective image, it is harmful (e.g. see Figure 1).

### v2

# OmniGuard: An Efficient Approach for AI Safety Moderation Across Languages and Modalities

@@ Abstract
The emerging capabilities of large language models (LLMs) have sparked concerns about their immediate potential for harmful misuse. The core approach to mitigate these concerns is the detection of harmful queries to the model. Current detection approaches are fallible, and are particularly susceptible to attacks that exploit mismatched generalization of model capabilities (e.g., prompts in low-resource languages or prompts provided in non-text modalities such as image and audio). To tackle this challenge, we propose OmniGuard, an approach for detecting harmful prompts across languages and modalities. Our approach (i) identifies internal representations of an LLM/MLLM that are aligned across languages or modalities and then (ii) uses them to build a language-agnostic or modality-agnostic classifier for detecting harmful prompts. OmniGuard improves harmful prompt classification accuracy by 11.57% over the strongest baseline in a multilingual setting, by 20.44% for image-based prompts, and sets a new SOTA for audio-based prompts. By repurposing embeddings computed during generation, OmniGuard is also very efficient ($\approx\!120\times$ faster than the next fastest baseline). Code and data are available at https://github.com/vsahil/OmniGuard.
@@
for AI Safety Moderation Across Languages and Modalities
@@ 1 Introduction
The rapid rise of capabilities in large language models (LLMs) has created an urgent need for safeguards to prevent their immediate harmful misuse as they are deployed to human users en masse bommasani2022-opportunities-risks. Moreover, these safeguards are critical for defending against future potential harms from LLMs bengio2024managing. Standard safeguard approaches broadly include approaches such as safety training using reinforcement learning from human feedback (instruct-gpt; leike2018-scalable-agent-alignment) or using pre-trained guard models that classify the safety of an input prompt  (openai2025moderation; inan2023-llama-guardrails; Han2024-wildguardmix-dataset).
With these safeguards in place, harmful prompts in high-resource languages, e.g., English, are successfully detected. However, harmful prompts in low-resource languages can often bypass these safeguards (deng2024-multilingual-attack-firstpaper; yong2024-low-resource-languages-jailbreak-gpt4; dan-roth-paper-guard-llms-on-multilingual-data), i.e., jailbreaking the LLM. Modern LLMs are vulnerable to attacks not only from low-resource natural languages, but also from artificial cipher languages, e.g., base64 or caesar encoding of English prompts (jacob-steinhardt-why-jailbroken; yuan2024-gpt-cipher-attack). This phenomenon also extends beyond text to jailbreaking multimodal LLMs (MLLMs) using modalities such as images (figstep-dataset-image-jailbreak; mm-safetybench-image-jailbreak) or audio (achilles-dataset-audio-jailbreaks).
jacob-steinhardt-why-jailbroken argue that these attacks are successful due to mismatched generalization, a scenario in which the model’s safety training does not generalize to other settings, but general performance does. This may happen because pretraining data often includes more diverse data than that available for safety finetuning (ghosh2024aegis2). In this work, we defend against attacks that exploit the mismatched generalization of the safety training of LLMs and MLLMs. Specifically, we defend against attacks that utilize low-resource languages, both natural and cipher languages, as well as attacks employing other modalities, such as images and audio.
@@
In contrast to popular guard models such as LlamaGuard (inan2023-llama-guardrails), AegisGuard (ghosh2024-aegis-safety-guard), or WildGuard (Han2024-wildguardmix-dataset), OmniGuard does not require training a separate LLM specifically to detect harmfulness. By building a classifier that uses the internal representations of the main LLM or MLLM, OmniGuard avoids the overhead of passing the prompt through a separate guard model, making it very efficient.
@@ 2 Methodology
OmniGuard seeks to robustly detect harmful prompts, regardless of their language or modality. We first leverage the tendency of LLMs and MLLMs to create universal representations that are similar across languages (wendler2024-convert-things-to-english; zhao2024-convert-things-to-english) and across modalities (wu2024-semantic-hub-hypothesis; zhuang2024vector) in Section˜2.1, and then use them to train harmfulness classifiers that robustly detect harmful inputs in Section˜2.2.
@@ 2.1 Finding language-agnostic representations in an LLM
This procedure can be generalized to new different modalities rather than different languages by changing which embeddings are being used. For example, to determine if internal representations of an MLLM are aligned across modalities, we replace embeddings for a translated piece of text with embeddings from a different modality (e.g. a text caption and its corresponding image, or a text transcription and its corresponding audio clip). See experimental details in Section˜3.
@@ 3 Experimental Setup
Table˜1 and Table˜2 give details on all the models and datasets for this section.
@@
General Flores200 nllbteam2022-languageleftbehindscaling Muennighoff/flores200 997
MM-Vet v2 yu2024-mmvetv2-dataset whyu/$m$-vet-v2 517
SST-2 (socher-etal-SSt-5-dataset) stanfordnlp/sst2 1000
Text Aegis AI Content Safety Dataset ghosh2024aegis2 nvidia/Aegis-AI-Content-Safety-Dataset-1.0 10,800
MultiJail deng2024-multilingual-attack-firstpaper DAMO-NLP-SG/MultiJail 315
Xsafety wang2024-multilingualsafety ToxicityPrompts/XSafety 28,000
RTP-LX rtplx-dataset ToxicityPrompts/RTP-LX 30,300
AyaRedTeaming aya-redteaming-dataset CohereLabs/aya_redteaming 2662
Thai Toxicity tweets thai-toxicity-tweets tmu-nlp/thai_toxicity_tweet 3,300
Ukr Toxicity ukr-toxicity-dataset ukr-detect/ukr-toxicity-dataset 5,000
HarmBench (HB) (mazeika2024-harmbench) walledai/HarmBench 400
Forbidden Questions (FQ) (forbidden-questions-dataset) TrustAIRLab/forbidden_question_set 390
Simple Safety Tests (simple-safety-tests) walledai/SimpleSafetyTests 100
SaladBench (SaladB) (salad-bench-dataset) walledai/SaladBench 26,500
Toxicity Jigsaw (TJS) toxicity-jigsaw Arsive/toxicity_classification_jigsaw 26,000
Toxic Text (toxic-text-dataset) nicholasKluge/toxic-text 41,800
AdvBench (advbench-dataset) walledai/AdvBench 520
CodeAttack (ren-2024-codeattack-jailbreak) https://github.com/AI45Lab/CodeAttack 3120
Vision JailBreakV-28K (luo2024jailbreakv28k) JailbreakV-28K/JailBreakV-28k 8,000
VLSafe (chen2024-vlsafe-dataset) YangyiYY/LVLM_NLF 1,110
FigStep (figstep-dataset-image-jailbreak) https://github.com/wangyu-ovo/MML 500
MML SafeBench (mml-safebench-image-jailbreak) https://github.com/wangyu-ovo/MML 2,510
Hades (hades-dataset-image-jailbreak) Monosail/HADES 750
SafeBench (another-safebench-dataset-image-jailbreaks) Zonghao2025/safebench 2,300
MM SafetyBench ($m$-safetybench-image-jailbreak) PKU-Alignment/$M$-SafetyBench 1680
RedTeamVLM (li2024-RedTeamVLM-dataset) MMInstruction/RedTeamingVLM 200
VLSBench (vlsbench-dataset-image-jailbreak) Foreshhh/vlsbench 2,240
Audio VoiceBench (Alpacaeval) advbench-audio-jailbreaks hlt-lab/voicebench 636
AIAH (achilles-dataset-audio-jailbreaks) https://github.com/YangHao97/RedteamAudioLMMs 350
@@ 3.1 Selecting universal layers via the U-Score
To select language-agnostic layers, we use a dataset of translated sentence pairs spanning various languages. Specifically, we use sentences in 53 natural languages from the Flores200 dataset and additionally translate the sentences into 20 cipher languages (using encodings such as Caesar shifts, base64, hexadecimal); see a full list in Appendix˜A. We extract embeddings from each layer of Llama3.3-70B-Instruct for the sentences in all 73 languages and use them to compute the U-Score (averaged over languages). Fig.˜2 shows the U-Score as a function of layer depth. For natural languages (blue curve), the U-Score peaks in the middle layers of the model, with the highest values in layer 57 (out of 81 layers). For cipher languages (red curve), the U-Score is much lower than for natural languages, suggesting the model fails to represent semantic similarity in these languages (see analysis in Section˜5).
@@
To select layers aligned between images and captions, we use the MM-Vet v2 dataset, a popular dataset for MLLM evaluation containing 517 examples, each consisting of a text question paired with one or more images. We generate captions for each image using a captioning model (Molmo-7B) and then extract embeddings for each image and its corresponding caption using an MLLM (also Molmo-7B) and use them to compute the U-Score, which peaks in layer 22 (out of 28 layers; see Fig.˜2 green curve).
To select layers aligned between text and audio, we use the audio version of the Alpacaeval dataset from VoiceBench, a dataset of 636 audio-transcript pairs. We extract embeddings from each layer of an MLLM (LLaMA-Omni 8B) and use them to compute the U-Score, which peaks at layer 21 (out of 32 layers; see Fig.˜2 purple curve).
@@ 3.2.1 Setup for multilingual text attacks
[…]
@@
We compare to many popular guard models (see Table˜2) middle row. Notably, DuoGuard and PolyGuard were trained to detect harmful prompts across multiple languages. For a more direct comparison, we also compare to finetuned versions of DuoGuard and PolyGuard using the same 37 languages we use to train OmniGuard; Following the original PolyGuard paper, we finetuned these models using LoRA (lora-paper) for all linear layers with rank 8 and alpha 16 for one epoch with a learning rate of 2e42e-4.
@@
We evaluate on several common text attack benchmarks (see Table˜1). We additionally evaluate on three benchmarks from CodeAttacks that transform a harmful query as a list, a stack, or as a string in a Python program, obfuscating the harmfulness. For evaluation in this setup, we transform the harmful prompts from AdvBench and benign prompts from Toxicity Jigsaw datasets in the three code formats and subsample the Toxicity Jigsaw dataset to be of the same size as Advbench. Note that for this experiment, we only trained OmniGuard on the English subset of the training dataset.
Model name Citation HuggingFace ID Rough Parameter Count
General Llama3.3-70B-Instruct (grattafiori2024llama) meta-llama/Llama-3.3-70B-Instruct 70B
Molmo-7B (deitke2024-molmo-model) allenai/Molmo-7B-D-0924 7B
LLaMA-Omni 8B (fang2025-llama-omni-model) ICTNLP/Llama-3.1-8B-Omni 8B
Kokoro (hexgrad_kokoro_82m_2025) hexgrad/Kokoro-82M 82M
Text LlamaGuard 1 inan2023-llama-guardrails meta-llama/LlamaGuard-7b 7B
LlamaGuard 2 inan2023-llama-guardrails meta-llama/Meta-Llama-Guard-2-8B 8B
LlamaGuard 3 inan2023-llama-guardrails meta-llama/Llama-Guard-3-8B 8B
AegisGuard Permissive ghosh2024-aegis-safety-guard nvidia/Aegis-AI-Content-Safety-LlamaGuard-Permissive-1.0 7B
AegisGuard Defensive ghosh2024-aegis-safety-guard nvidia/Aegis-AI-Content-Safety-LlamaGuard-Defensive-1.0 7B
WildGuard Han2024-wildguardmix-dataset allenai/wildguard 7B
HarmBench (mistral) mazeika2024-harmbench cais/HarmBench-Mistral-7b-val-cls 7B
HarmBench (llama) mazeika2024-harmbench cais/HarmBench-Llama-2-13b-cls 13B
DuoGuard duo-guard-multilingual DuoGuard/DuoGuard-1B-Llama-3.2-transfer 1B
PolyGuard polyguard-model ToxicityPrompts/PolyGuard-Qwen 7B
Vision Llama Guard 3 Vision (chi2024-llamaguard3-vision) meta-llama/Llama-Guard-3-11B-Vision 11B
VLMGuard (du2024-vlmguard) --- 2.2M
LLavaGuard (helff2025-llavaguard) AIML-TUDA/LlavaGuard-7B-hf 7B
Table 2: Model and baseline details.
@@ 3.2.2 Setup for vision attacks
Following Section˜3.1, we build a classifier that takes as input embeddings from layer 22 of Molmo-7B. As training data, we use 2000 image-query pairs randomly sampled from the JailBreakV-28K dataset and 1024 image-query pairs sampled from the VLSafe dataset as the harmful datapoints and 517 image-query from the MM-Vet v2 dataset as the benign datapoints.
@@
We compare to guard models that take an image or image-text pair and output a binary harmfulness classification (see Table˜2 bottom row). We train VLMGuard on the same training data as OmniGuard.
@@
We evaluate detecting image/text attacks using several datasets (see Table˜1). FigStep and MML Safebench are typographic attacks that embed a harmful prompt in an image. MML Safebench further encrypts a harmful prompt in several variants, such as rotation, mirror images, word replacement, and with base64 encoding. Hades and Safebench consist of images and text queries where the text itself is harmful. MM-safetybench, RTVLM, and VLSBench consist of an image and a query where the text query is seemingly benign, but when combined with the respective image, it is harmful (e.g. see Figure˜1).